# Lessons from the TREC Plain Language Adaptation of Biomedical Abstracts (PLABA) track

### v1

@@ Abstract
{highlights}
We introduce the new shared task of plain language adaptation of biomedical abstracts.
Teams from twelve countries submitted runs with systems based on BERT, BART, T5, GPT, Llama, Gemini, and Mistral.
In manual judgments, top-performing models neared human levels of factual accuracy, but not simplicity or brevity.
Automatic, reference-based metrics generally did not correlate well with manual judgments.
@@ 1 Introduction
The inability of patients to fully understand available information about their health has a significant impact on outcomes (Stableford and Mettger, 2007; Berkman et al., 2011; King, 2010; Kindig et al., 2004). While more information than ever is available on the web, the “language barrier” of professional-facing content, such as literature, can make it counterproductive (Rosenberg et al., 2017; White and Horvitz, 2009). Though many consumer-facing knowledge bases exist, these are cumbersome and labor-intensive to update and thus typically do not include the latest medical knowledge from the literature. This is especially important for consumers seeking out the latest information about emerging health issues (SeyyedHosseini et al., 2022).
The many recent, successful applications of Deep Learning to the field of Natural Language Processing suggest the possibility of automatic adaptation of text intended for a professional audience to text that can be read and understood by a consumer audience. Though Text Simplification has been widely studied in the open domain (Al-Thanyyan and Azmi, 2021), the unique lexicon of biomedical literature makes it a distinct sublanguage (Friedman et al., 2002). In this respect, the task overlaps with Machine Translation (Stahlberg, 2020). Recent examples of applying Deep Learning to Text Simplification specifically in the biomedical domain include Van den Bercken et al. (2019), Devaraj et al. (2021), and Flores et al. (2023). However, the existing body of research evaluates systems with disparate datasets that are mined from imperfect sources, making progress difficult to measure.
@@ 2 Related Work
In addition to being focused solely on the biomedical domain, PLABA differentiates from SimpleText in the extent of manual evaluation. Evaluation of SimpleText Task 3 largely focuses on automatic, reference-based metrics, such as BLEU (Papineni et al., 2002) and SARI (Xu et al., 2016), and count-based readability measures, such as FKGL (Kincaid et al., 1975). The 2024 offering additionally included analysis of spurious content (i.e. hallucinations) via alignment against sources. Extensive manual evaluation of a subset of outputs was performed for SimpleText 2022’s Task 3, but only for the Computer Science corpus. Though manual evaluation was performed for SimpleText 2024’s Task 3, reported results were limited to overall observations based on a small sample of outputs. PLABA’s Task 1 adds to this work by providing in-depth manual evaluation for a purely biomedical test corpus, judging outputs for every sentence of 40 abstracts in 2023 and 400 abstracts in 2024.
@@ 3.1.1 Task Definition
For a set of 40 consumer health questions, teams are presented with ten abstracts relevant to each question, each one split into sentences. The task is to create a sentence-aligned plain language adaptation of each abstract (Fig. 1). Being sentence-aligned means that output must be provided for each source sentence, but the entire rewritten abstract is also expected to read fluently as one document (e.g. not repeatedly explain expert terms). As in the PLABA dataset, sentences may be split, meaning output for a single source sentence can have multiple sentences, and this is in fact encouraged when source sentences are long and complex. Sentences may also be dropped if they are deemed not relevant to a consumer’s understanding of the abstract. However, sentences may not be merged (i.e. providing a single output sentence that spans the semantic content of multiple source sentences), to ensure there is a direct and consistent mapping from source to output for evaluation purposes.
@@ 3.2 Task 2: Identifying and Replacing Complex Terms
The initial offering of Task 1 (end-to-end rewriting of abstracts) at TREC 2023 revealed that language models on the order of billions of parameters were necessary for competitive performance. This could have limited participation in the first offering, as expensive, state-of-the-art GPUs and servers are required to fine-tune such models. Zero-shot configurations, on the other hand, required recent commercial models, which were also associated with significant costs. For the second year of PLABA at TREC, we thus designed a new sub-task around identifying and replacing terms, rather than complete rewriting. One goal of this task was partly to lower the barrier to entry by including shorter generations and classification tasks. Another was to generate finer-grained feedback for the lexical aspect of plain language adaptation.
@@ 3.2.1 Task Definition
    Task 2A - Identification: The first sub-task was to identify difficult terms. Given the text of an entire abstract, systems should return a list of unique substrings representing words or phrases that a consumer would not understand.
@@
    Task 2B - Classification: Given identified expert terms, the second sub-task was to classify how the terms should be replaced. As there may be multiple valid ways a human writer could handle replacing a given term, this is framed as a multi-label problem (that is, a binary classification problem for each of the five labels).
@@
    Task 2C - Generation: The third sub-task was the generate simplifications (replacement terms for substitute and generalize or additional text for explain and exemplify) given an abstract and the expert terms identified and classified within it. Note that if a term was classified only as omit in Task 2B, it would not have any generation for Task 2C.
@@ 4.1.2 Manual Evaluation
We contracted four science writing experts to perform the annotations. Annotators were provided with rubrics for 3-point likert scale (-1, 0, 1) judgments for each axis and trained via a live video session. For final results, averages of likert scale values were linearly interpolated from (-1, 1) to (0, 100) for easier interpretation. As a pilot, all four annotated the first abstract, after which significant differences were discussed and questions clarified during a followup video session. During full evaluation of systems, a subset was double-annotated to compute inter-annotator agreement (see Appendix G).
@@ 4.1.3 Automatic Sentence Alignment
Though PLABA is a sentence-level task, we anticipated that the importance of context would make document-level systems attractive. However, such systems typically do not assure sentence-alignment, which we require for evaluation. To give teams this option without the added burden of aligning output sentences, we provided an automatic alignment pipeline for document-level submissions. This pipeline was based on Vecalign (Thompson and Koehn, 2019), which was developed for bilingual alignment for Machine Translation tasks but worked well in this setting. Vecalign implements efficient dynamic programming using embeddings to score similarity. Additionally, embeddings of overlapping blocks of sentences are used to allow many-to-many alignment. As in Thompson and Koehn (2019), we use LASER (Artetxe and Schwenk, 2019) for embeddings. Initial sentence-segmentation of document-level system outputs was performed using Stanza (Qi et al., 2020). Finally, since Vecalign allows insertions in one-to-many mode, but the PLABA task does not, we also developed an assignment algorithm to handle these cases. Additional details can be found in Appendix A.
@@ 4.2 Task 1 at TREC 2024
Instead of fluency, we introduce brevity (BRV), which captures whether systems produce output that conveys information as concisely as possible (taking into account how much information should be conveyed). This aspect of outputs had been conflated with SEN (sentence simplicity) in the initial offering, which caused confusion among annotators and contributed to poor inter-annotator agreement for this metric (see Appendix G). We thus deemed it better to provide it as a separate axis for the second offering.
@@ 5.1.1 Teams and Systems
Team submissions are detailed in Appendix C. Each team was initially allowed three submissions, which they could rank in order of priority for manual evaluation. We also asked each team to provide short description of the strategy of each submission.
5.1.2 Baseline Systems
For the inaugural offering of PLABA, we did not know how many submissions to expect, and thus developed a handful of our own baseline systems. These explored various prompting strategies as well as both zero-shot proprietary models and fine-tuned open-source models.
  • 1.
    PLABA_1: Sentence-level adaptation via zero-shot auto-regressive prompting of text-davinci-003, a variant of OpenAI’s proprietary GPT 3.5 base model with instruction tuning via Reinforcement Learning from Human Feedback (Ouyang et al., 2022). This involved progressively building prompts for each sentence of the abstract, using prior model outputs as in-context examples. This inherently assigned outputs to individual source sentences, but also encouraged use of context when adapting each sentence. Note that, for a given source sentence, the model may return multiple sentences, essentially performing a split operation. However, these can still be attributed to a single source sentence, so alignment is not required. Prompt details are provided in  Appendix D.
  • 2.
  • 3.
  • 4.
  • 5.
    PLABA_5: The same model as PLABA_4, but fine-tuned without SCER.
@@ 5.1.4 Manual Evaluation Results
As further resources were available following the first round of manual evaluation, we invited teams to submit an additional run for a second round of manual evaluation. As only one team (BeeManc) did so, we manually evaluated the second priority runs for the other teams, except PT3M, who had only submitted one run. Resources also allowed us to include two additional baselines, both fine-tuned (PLABA_4 and PLABA_5), and to include the manually written reference adaptations for manual evaluation, creating a human performance baseline. Finally, since different annotators were used versus the first round, we included the top-performing system (PLABA_1) for re-evaluation as a point of reference. Note, however, that these results are not directly comparable to those in round 1 because of the change in annotators and timing (after the conference rather than before). Results of both rounds of manual evaluation are shown in Table 3 and visualized in Figure 2(a). During the second round of evaluation, resources allowed for giving annotators overlapping assignments for three abstracts in order to estimate inter-annotator agreement (see Appendix G).
@@ 5.1.5 Analysis of Metrics
To assess the utility of automated metrics, we compute their correlation with manual judgments. Correlations of all computed metrics are shown in Figure 3(a). Scatterplots with regression lines for each metric are shown in Appendix H.
@@ 5.3 Task 2 at TREC 2024
Six teams participated in Task 2 at TREC 2024, with nine total submissions. Submissions were due September 20th, and preliminary results were announced November 10th, 2024. Team submissions are detailed in Appendix F. Each team was allowed three submissions, which they could rank in order of priority for manual evaluation. Tasks 2B and 2C were optional, though participation in 2C required participation in 2B. We also asked each team to provide a short description of each submission, as well as any base LLMs and training data used.
5.3.1 Baseline Systems
We developed baseline systems for the identification (Task 2A) and classification (Task 2B) sub-tasks within Task 2. This included three baselines for Task 2A:
  • 1.
  • 2.
  • 3.
  • 1.
  • 2.
@@ 5.3.2 Evaluation Results
All results for Task 2 at TREC 2024, including automatic metrics for Tasks 2A and 2B and manual judgments for Task 2C, are shown in Table 5. As for Task 1, we compute correlations among each pair of manual evaluation axes for Task 2C (Fig. 3(c)). Again, accuracy is strongly related to completeness; however, in this case, brevity and completeness exhibit a negative correlation, illustrating the tradeoff between outputs being complete and brief. Pairwise scatterplots and correlations broken down by replacement type are shown in Appendix J.
@@ 6.3 Factuality and Hallucinations
Still, this shared task also revealed that, even when generation is narrowly focused on one source line, large, convincing hallucinations can still occur (as described in Appendix I). These may be all the more insidious if users learn to trust systems that are largely accurate. Health information provided to consumers can be highly actionable, and there is thus a high potential for harm even in rare edge cases. We must thus be vigilant in evaluating and deploying systems that provide such information. Future work should investigate more rigorous ways to detect and mitigate hallucinations and automatically assess factuality to ensure these kinds of errors do not go undetected.
@@ 8 Acknowledgements
The entire pipeline for aligning sentences of document-level submissions to reference sentences is depicted in Figure 1. One-to-many alignments from Vecalign allow insertions versus the source; i.e., system blocks that are marked as aligned between two source sentences, or at the beginning or end of the source. However, the PLABA task does not allow insertions, and all outputs must be assigned to a source sentence for evaluation. Any system blocks marked as insertions by the alignment process thus need to be joined to either the previous or following system block. We choose, arbitrarily, to join them to the following system block, unless the inserted block is terminal (marked as inserted after the last source sentence, with no system blocks following), in which case we join it with the previous system block. The algorithm is depicted in Algorithm A.1.
@@
Team submissions to Task 1 at TREC 2023, and their descriptions, are shown in Table 1.
@@
Appendix Appendix D Prompting strategy for PLABA_1 baseline
The sentence-wise auto-regressive zero-shot prompting strategy for the baseline system PLABA_1 is shown in Figure 2. An initial prompt contains a general instruction (“Rewrite for a lay audience:”), the first sentence of the abstract with a label (“Original:”), and a dangling label (“Simple:”). The model’s response to this prompt is then included in the next prompt, which has a dangling label for the second simple sentence. The process repeats for each sentence in the original abstract, storing model responses for each.
Figure 2: The sentence-wise auto-regressive zero-shot prompting strategy for the baseline system PLABA_1.
Appendix Appendix E System Descriptions for Task 1 at TREC 2024
@@
Appendix Appendix F System Descriptions for Task 2 at TREC 2024
@@
Appendix Appendix G Inter-annotator Agreement for Manual Evaluation of Task 1
@@
Appendix Appendix H Scatterplots of Automatic Metrics Versus Manual Judgments of Task 1 at TREC 2023
@@
Appendix Appendix I Hallucinations from Task 1 at TREC 2023
@@
Appendix Appendix J Scatterplots of Manual Evaluation Axes for Task 2 by Replacement Type

### v2

@@ 1 Introduction
The inability of patients to understand available information about their health significantly impacts outcomes (Stableford and Mettger, 2007; Berkman et al., 2011; King, 2010; Kindig et al., 2004). While more information than ever is available on the web, the “language barrier” of professional-facing content, such as literature, can make it counterproductive (Rosenberg et al., 2017; White and Horvitz, 2009). Though many consumer-facing knowledge bases exist, these are cumbersome and labor-intensive to update and thus typically do not include the latest medical knowledge from the literature. This is especially important for consumers seeking out the latest information about emerging health issues (SeyyedHosseini et al., 2022).
The many recent successes of Artificial Intelligence suggest the possibility of automatically adapting text intended for a professional audience to a consumer audience. Though Text Simplification has been widely studied in the open domain (Al-Thanyyan and Azmi, 2021), the unique lexicon of biomedical literature makes it a distinct sublanguage (Friedman et al., 2002). In this respect, the task overlaps with Machine Translation (Stahlberg, 2020). Recent examples of applying Deep Learning to Text Simplification specifically in the biomedical domain include Van den Bercken et al. (2019), Devaraj et al. (2021), and Flores et al. (2023). However, the existing body of research evaluates systems with disparate datasets that are mined from imperfect sources, making progress difficult to measure.
@@
Statement of Significance
Problem
Patients and caretakers cannot keep up with biomedical research due to technical language and jargon.
What is Already Known
Advanced Artificial Intelligence systems have sought to simplify professional text for a general audience.
What This Paper Adds
We stimulate methods research by providing a shared task and provide extensive manual review of outputs by biomedical experts to assess accuracy.
Who would benefit from the knowledge in this paper
Researchers using Artificial Intelligence to generate public-facing biomedical text output; clinicians who need to explain current research to patients.
@@ 2 Related Work
In addition to focusing on the biomedical domain, PLABA differentiates from SimpleText in the extent of manual evaluation. Evaluation of SimpleText Task 3 largely focuses on automatic, reference-based metrics, such as BLEU (Papineni et al., 2002) and SARI (Xu et al., 2016), and count-based readability measures, such as FKGL (Kincaid et al., 1975). The 2024 offering additionally included analysis of spurious content (i.e. hallucinations) via alignment against sources. Extensive manual evaluation of a subset of outputs was performed for SimpleText 2022’s Task 3, but only for the Computer Science corpus. Though manual evaluation was performed for SimpleText 2024’s Task 3, reported results were limited to overall observations based on a small sample of outputs. PLABA’s Task 1 adds to this work by providing in-depth manual evaluation for a purely biomedical test corpus, judging outputs for every sentence of 40 abstracts in 2023 and 400 abstracts in 2024.
@@ 3.1.1 Task Definition
For a set of 40 consumer health questions, teams are presented with ten abstracts relevant to each question, each one split into sentences. The task is to create a sentence-aligned plain language adaptation of each abstract (Fig. 1). Being sentence-aligned means that output must be provided for each source sentence, but the entire rewritten abstract is also expected to read fluently as one document (e.g. not repeatedly explain expert terms). As in the PLABA dataset, sentences may be split (multiple output sentences for a single source sentence), and this is in fact encouraged when source sentences are long and complex. Sentences may also be dropped if they are deemed not relevant to a consumer’s understanding of the abstract. However, sentences may not be merged (i.e. providing a single output sentence that spans the semantic content of multiple source sentences), to ensure there is a direct and consistent mapping from source to output for evaluation purposes.
@@ 3.2 Task 2: Identifying and Replacing Complex Terms
For the second year of PLABA at TREC, we designed a new sub-task around identifying and replacing terms, rather than complete rewriting. One goal of this task was partly to lower the barrier to entry by including shorter generations and classification tasks, as the initial offering of Task 1 revealed that language models on the order of billions of parameters were necessary for competitive performance. Another goal was to generate finer-grained feedback for the lexical aspect of plain language adaptation.
@@ 3.2.1 Task Definition
    Task 2A - Identification: Given the text of an entire abstract, systems should return a list of unique substrings representing words or phrases that a consumer would not understand.
@@
    Task 2B - Classification: Given identified expert terms, classify how the terms should be replaced. As there may be multiple valid ways a human writer could handle replacing a given term, this is framed as a multi-label problem (that is, a binary classification problem for each of the five labels).
@@
    Task 2C - Generation: Generate simplifications (replacement terms for substitute and generalize or additional text for explain and exemplify) given an abstract and the expert terms identified and classified within it. Note that if a term was classified only as omit in Task 2B, it would not have any generation for Task 2C.
@@ 4.1.2 Manual Evaluation
We contracted four science writing experts to perform the annotations. Annotators were provided with rubrics for 3-point likert scale (-1, 0, 1) judgments for each axis and trained via a live video session. For final results, averages of likert scale values were linearly interpolated from (-1, 1) to (0, 100) for easier interpretation. As a pilot, all four annotated the first abstract, after which significant differences were discussed and questions clarified during a followup video session. During full evaluation of systems, a subset was double-annotated to compute inter-annotator agreement (see Appendix F).
@@ 4.1.3 Automatic Sentence Alignment
Though PLABA is a sentence-level task, we anticipated that the importance of context would make document-level systems attractive. However, such systems typically do not assure sentence-alignment, which we require for evaluation. To give teams this option without the added burden of aligning output sentences, we provided an automatic alignment pipeline for document-level submissions. Additional details can be found in Appendix A.
@@ 4.2 Task 1 at TREC 2024
Instead of fluency, we introduce brevity (BRV), which captures whether systems produce output that conveys information as concisely as possible (taking into account how much information should be conveyed). This aspect of outputs had been conflated with SEN (sentence simplicity) in the initial offering, which caused confusion among annotators and contributed to poor inter-annotator agreement for this metric (see Appendix F). We thus deemed it better to provide it as a separate axis for the second offering.
@@ 5.1.1 Teams and Systems
Team submissions and baselines are detailed in Appendix C. Each team was initially allowed three submissions, which they could rank in order of priority for manual evaluation. We also asked each team to provide short description of the strategy of each submission.
@@ 5.1.3 Manual Evaluation Results
As further resources were available following the first round of manual evaluation, we invited teams to submit an additional run for a second round of manual evaluation. As only one team (BeeManc) did so, we manually evaluated the second priority runs for the other teams, except PT3M, who had only submitted one run. Resources also allowed us to include two additional baselines, both fine-tuned (PLABA_4 and PLABA_5), and to include the manually written reference adaptations for manual evaluation, creating a human performance baseline. Finally, since different annotators were used versus the first round, we included the top-performing system (PLABA_1) for re-evaluation as a point of reference. Note, however, that these results are not directly comparable to those in round 1 because of the change in annotators and timing (after the conference rather than before). Results of both rounds of manual evaluation are shown in Table 3 and visualized in Figure 2(a). During the second round of evaluation, resources allowed for giving annotators overlapping assignments for three abstracts in order to estimate inter-annotator agreement (see Appendix F).
@@ 5.1.4 Analysis of Metrics
To assess the utility of automated metrics, we compute their correlation with manual judgments. Correlations of all computed metrics are shown in Figure 3(a). Scatterplots with regression lines for each metric are shown in Appendix G.
@@ 5.3 Task 2 at TREC 2024
Six teams participated in Task 2 at TREC 2024, with nine total submissions. Submissions were due September 20th, and preliminary results were announced November 10th, 2024. Team submissions and baseline systems are detailed in Appendix E. Each team was allowed three submissions, which they could rank in order of priority for manual evaluation. Tasks 2B and 2C were optional, though participation in 2C required participation in 2B. We also asked each team to provide a short description of each submission, as well as any base LLMs and training data used.
@@ 5.3.1 Evaluation Results
All results for Task 2 at TREC 2024, including automatic metrics for Tasks 2A and 2B and manual judgments for Task 2C, are shown in Table 5. As for Task 1, we compute correlations among each pair of manual evaluation axes for Task 2C (Fig. 3(c)). Again, accuracy is strongly related to completeness; however, in this case, brevity and completeness exhibit a negative correlation, illustrating the tradeoff between outputs being complete and brief. Pairwise scatterplots and correlations broken down by replacement type are shown in Appendix I.
@@ 6.3 Factuality and Hallucinations
Still, this shared task also revealed that, even when generation is narrowly focused on one source line, large, convincing hallucinations can still occur (as described in Appendix H). These may be all the more insidious if users learn to trust systems that are largely accurate. Health information provided to consumers can be highly actionable, and there is thus a high potential for harm even in rare edge cases. We must thus be vigilant in evaluating and deploying systems that provide such information. Future work should investigate more rigorous ways to detect and mitigate hallucinations and automatically assess factuality to ensure these kinds of errors do not go undetected.
@@ 8 Acknowledgements
The entire pipeline for aligning sentences of document-level submissions to reference sentences is depicted in Figure 1. This pipeline was based on Vecalign (Thompson and Koehn, 2019), which was developed for bilingual alignment for Machine Translation tasks but worked well in this setting. Vecalign implements efficient dynamic programming using embeddings to score similarity. Additionally, embeddings of overlapping blocks of sentences are used to allow many-to-many alignment. As in Thompson and Koehn (2019), we use LASER (Artetxe and Schwenk, 2019) for embeddings. Initial sentence-segmentation of document-level system outputs was performed using Stanza (Qi et al., 2020). Finally, since Vecalign allows insertions in one-to-many mode, but the PLABA task does not, we also developed an assignment algorithm to handle these cases. One-to-many alignments from Vecalign allow insertions versus the source; i.e., system blocks that are marked as aligned between two source sentences, or at the beginning or end of the source. However, the PLABA task does not allow insertions, and all outputs must be assigned to a source sentence for evaluation. Any system blocks marked as insertions by the alignment process thus need to be joined to either the previous or following system block. We choose, arbitrarily, to join them to the following system block, unless the inserted block is terminal (marked as inserted after the last source sentence, with no system blocks following), in which case we join it with the previous system block. The algorithm is depicted in Algorithm A.1.
@@
Team submissions to Task 1 at TREC 2023, and their descriptions, are shown in Table 1. For the inaugural offering of PLABA, we did not know how many submissions to expect, and thus developed a handful of our own baseline systems. These explored various prompting strategies as well as both zero-shot proprietary models and fine-tuned open-source models.
  • 1.
    PLABA_1: Sentence-level adaptation via zero-shot auto-regressive prompting of text-davinci-003, a variant of OpenAI’s proprietary GPT 3.5 base model with instruction tuning via Reinforcement Learning from Human Feedback (Ouyang et al., 2022). This involved progressively building prompts for each sentence of the abstract, using prior model outputs as in-context examples (Fig. 2). This inherently assigned outputs to individual source sentences, but also encouraged use of context when adapting each sentence. Note that, for a given source sentence, the model may return multiple sentences, essentially performing a split operation. However, these can still be attributed to a single source sentence, so alignment is not required.
  • 2.
  • 3.
  • 4.
  • 5.
    PLABA_5: The same model as PLABA_4, but fine-tuned without SCER.
@@
Figure 2: The sentence-wise auto-regressive zero-shot prompting strategy for the baseline system PLABA_1. An initial prompt contains a general instruction (“Rewrite for a lay audience:”), the first sentence of the abstract with a label (“Original:”), and a dangling label (“Simple:”). The model’s response to this prompt is then included in the next prompt, which has a dangling label for the second simple sentence. The process repeats for each sentence in the original abstract, storing model responses for each.
Appendix Appendix D System Descriptions for Task 1 at TREC 2024
@@
Appendix Appendix E System Descriptions for Task 2 at TREC 2024
@@
We also developed baseline systems for the identification (Task 2A) and classification (Task 2B) sub-tasks within Task 2. This included three baselines for Task 2A:
  • 1.
  • 2.
  • 3.
  • 1.
  • 2.
@@
Appendix Appendix F Inter-annotator Agreement for Manual Evaluation of Task 1
@@
Appendix Appendix G Scatterplots of Automatic Metrics Versus Manual Judgments of Task 1 at TREC 2023
@@
Appendix Appendix H Hallucinations from Task 1 at TREC 2023
@@
Appendix Appendix I Scatterplots of Manual Evaluation Axes for Task 2 by Replacement Type